OpenStudy (anonymous):

How do I solve this? 7+(5-2)-(| -12| -4)÷(8-6) and what kind of problem is this?

OpenStudy (anonymous):

Resolve the parens/brackets/absolute values first, starting with the innermost ones. Then perform the remaining operations in order of precedence. Exponentiate before multiplication/division before addition/subtraction and always from left to right.

OpenStudy (anonymous):

ok thanks.. also would this be the way to simplify the problem too? because actually that is what I need to do

OpenStudy (anonymous):

Yes, there is nothing to 'solve' here. You have an expression that can be simplified. What I described was how to simplify it. I assumed you meant simplify, since few people actually understand what it is to solve something, and often use the two interchangeably.

OpenStudy (anonymous):

ok thank you much

OpenStudy (anonymous):

Of course! If you like I can assist while you work through it. In case you run into trouble..

OpenStudy (anonymous):

7+(5-2)-( |-12|-4)÷(8-6) 7+3-(12-4)/(8-6) 7+3=10-8/4=2 is this right?

OpenStudy (anonymous):

u still with me polpak?

OpenStudy (anonymous):

I don't like the notation you used on the last line there, 7+3 is 10. 8/4 is 2. 7+3-8/4 = 10-2

OpenStudy (anonymous):

If that is what you mean, then that is correct.

OpenStudy (anonymous):

ok but the answer is different than my book.it says the answer is 6,but i can never get that answer.also when i did a yahoo ask question search they also got 6.i just can't seem to figure out how

OpenStudy (anonymous):

ok think i got it now... 7 + (5 - 2) - ( |-12| - 4) ÷ (8 - 6) = 7 + 3 - 8 ÷ 2 = (7 + 3) - (8 ÷ 2) = 10 - 4 = 6

OpenStudy (anonymous):

\(7+(5-2)-(| -12| -4)÷(8-6)\) \[= 7 + 3 - (12 - 4) \div 2\]\[=10-8\div 2\]\[=10 - 4\]\[=6\]

OpenStudy (anonymous):

thank u so much 4 helping me.... :)

OpenStudy (anonymous):

No problem =)